OpenStudy (jkbo):

What is the solution to the proportion? 3y - 8 over 12 = y/5

OpenStudy (anonymous):

\[\Large \frac{3y-8}{12} = \frac{y}{5}\] You can solve this one of two different ways. 1. Use cross multiplication. 2. Multiply both sides of the equation by 1 in the form of a fraction so that the denominators are the same. Then solve the equation using the numerators.

OpenStudy (jkbo):

thanks